OpenStudy (anonymous):

candice is the bst hitter on her high school softball team. She gets on base safely 60% of the time. What is the probability that her next at bat will result in her getting on base

OpenStudy (anonymous):

well she gets on the base 60% of the time. convert 60% into the probability

OpenStudy (anonymous):

it can be written as percentage - in which the answer is already given or it can be given as a fraction

OpenStudy (anonymous):

A. 3/5 B. 1/6 C.1/.2 D. 2/5

OpenStudy (anonymous):

@touseii45

OpenStudy (anonymous):

whats another way of writing 60% in fractions?

OpenStudy (anonymous):

if you get a 60% grade on your homework, what does it mean?

OpenStudy (anonymous):

1/6?

OpenStudy (anonymous):

not neccessarily, 1/6 would mean you got 1 out of 6 questions correct.

OpenStudy (anonymous):

60% would mean you got 60/100 questions correct

OpenStudy (anonymous):

what if your homework has only 10 questions, how many did you get correct if you got 60% on it?

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

yes its 3/5 because 3/5= 6/10= 60/100 = 60%

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

alternatively you can just plug 3 divided by 5 to see if you get .60 as an answer.

OpenStudy (anonymous):

yea what they said lol!

OpenStudy (anonymous):

lml